OpenStudy (anonymous):

Find sin 2x, cos 2x, and tan 2x from the given information. tan x = (− 12/5), x in Quadrant II

myininaya (myininaya):

find hyp=sqrt(opp^2+adj^2)

myininaya (myininaya):

\[\sin(2x)=2\sin(x)\cos(x)=2 \frac{opp}{hyp} \frac{adj}{hyp}\]

myininaya (myininaya):